A cylinder of mass 5kg and radius 30cm, is free to rotate about its axis;it receives an angular impulse of 3kg-m2-s-1 initially followed by a similar impulse after every 4s. What is the angular speed of the cylinder after 30s of the initial impulse? The cylinder was at rest initially.
 

Grade:12th pass

1 Answers

Suman Panigrahi
6 years ago
The cylinder experiences 8 impulse of 3kg-m^2/s each i.e. 24kg-m^2/s . Since L=IW so Iw=24. I for cylinder =MR^2÷2.so I=0.45÷2 . Hence w=(24×2)÷0.45=106.7